OpenStudy (anonymous):

What's the range for f(x)=4/(x^2-3x)

OpenStudy (baldymcgee6):

http://gyazo.com/639abde5d6c93ade3cc255f688ffa849

OpenStudy (eyust707):

What are all the possible y values that could come from that function?

OpenStudy (anonymous):

or one can find the domain of inverse

OpenStudy (anonymous):

But can not find the inverse

OpenStudy (anonymous):

what @mukushla said. probably the best bet

OpenStudy (anonymous):

wrong function

OpenStudy (anonymous):

\[f(x)=\frac{4}{x^2-3x}\]

OpenStudy (anonymous):

if you want to use wolfram and not compute by hand, it is this one http://www.wolframalpha.com/input/?i=range+f%28x%29%3D \frac{4}{x^2-3x}

OpenStudy (anonymous):

which i guess requires copy and paste

OpenStudy (eyust707):

what about the inverse?

OpenStudy (anonymous):

or you can take \[y=\frac{4}{x^2-3x}\] and solve for \(x\) to see what you get, as @mukushla suggested that will give you the range

OpenStudy (eyust707):

ohhh

OpenStudy (anonymous):

@eyust707 mukushia means inverse function, not reciprocal in other words solve for \(x\)

OpenStudy (eyust707):

got it! thx for the explanation!

OpenStudy (anonymous):

yw, hope whoever asked also knows what to do

OpenStudy (anonymous):

I got it. I figured out the inverse function by using quadratic formula, and found out the range for it. Thanks everybody.

OpenStudy (anonymous):

very well :)

OpenStudy (anonymous):

that is the right thing to do you get a radical, make sure the discriminant is \(\geq 0\)